How many microstates are there for the 2p1 3p1 (one electron in 2p, and another in 3p orbitals) configuration?
Answer #1

1 Total spin of this arrangement S=&ms = 1 + 1 = 1 Spin multiplicity = (25+1) = (2x1 +1) = 5 3 Total angular momentum L={me =